Erika drives her car with an average velocity of 24 m/s toward the east.
a. How long will it take her to drive 560 km on a perfectly straight highway?
b. How much time would she save if she increased her average velocity to 26 m/s to the east?

Answer by mananth(16949)

1 m/s = 3.6 km/h
24m/s = 24*3.6
=86.4 km/h
time = distance/speed
560/86.4
= 6.48 hours
...
26m/s= 26*3.6=93.6 km/h
560/93.6
= 5.98 hours
...
0.5 hours saved
...
m.ananth@hotmail.ca